OpenStudy (anonymous):

Part 1: For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 10 3x + 4y = 25

OpenStudy (anonymous):

just a thought, algebra and system of algebraic equations is useful in the field of economics. .

OpenStudy (anonymous):

we want either x's or y's to be equal and opposite to cancel: so we try x's. (x-3) -6x -3y = -30 muliply across by -3 (x +2) 6x +8y= 50 multiply across by +2 result: -6x and 6x cancel -3y+8y=5y -30+50=20 that is: 5y=20 y=20/5 y=4 substitute y=4 in eq 1. 2x+y=10 2x+(4)=10 2x=10-4 2x=6 x=6/2 x=3 check in 2nd eq for x=3, y=4 3(3)+4(4)=25 9+16=25 25=25 Q.E.D.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@BPDlkeme , where is the real world scenario there?

OpenStudy (anonymous):

well if you re-arrange the equations (which are linear equations, i.e. equations of the first degree) then the solution x=3, y=4 represents where these two lines cross. A real world example using these types of equations could be taken from linear programming. Say for example you produce x widgets, obviously the more x widgets you produce, the higher the cost. if this was shown graphically (cost vs units produced) it could represent one line, and on the other you might have profit v sales (snother linear equation. Where the two lines intersect (x=3, y=4) would be your break-even point, i.e. costs=sales, anything above this would be profit.; HTH.